OpenStudy (anonymous):

a car traveling at 20 m s accelerates uniformly at 3 m/s^2 for 5s. calculate the distance the car travels while it is accelerating. i need how to work it out more than the answer

OpenStudy (anonymous):

Do you know the distance formula?

OpenStudy (anonymous):

s = d/t so d = s X t ?

OpenStudy (anonymous):

\(d=v_it + \frac{1}{2}at^2\), since a is uniform, you can use this formula. where \(v_i= 20 m/s\), \(a=3 m/s^2\) and \(t=5s\)